A survey on the background concentration of lead in soil was conducted in an Environmental Impact Assessment of a former battery plant. Preliminary results of six soil samples in the protected area give results in mg/kg: 28, 32, 21, 29, 25, 22. (a) Is there any value in this data set that can be discarded as an outlier at 90% confidence level? (b) Calculate CV, RSD, and 95% Confidence Interval (CI)?


1
Expert's answer
2021-09-07T23:28:00-0400

a) 28 + 32 +21 + 29 + 25+22 = 157

157/6 = 26.167 × 0.9 = 23.55

Hence, 21 and 22